Suppose the prices of one-year, two-year, and three-year zero coupon bonds each with a par value of $100 are $90, $80, and $70, respectively.

Compute the three spot rates.
Answer #1

The formula for the zero-coupon bond is -

Price = Par Value /(1+ Spot Rate )n

n = number of periods

Given that

Par Value of all three bonds = $100

1. 90 = 100/(1+s1)1

s1 = one year spot rate = 1/9 = 11.11%

2. 80 = 100/(1+s2)2

s2 = two year spot rate = 1.1180-1 = 11.80%

3. 70 = 100/(1+s3)3

s3 = three year spot rate = 1.2162-1 = 12.62%